Please help!!
Question:
A box contains 4 red and 7 green marbles. You reach in and remove 2 marbles all at once. What is the probability of picking one red marble, and one green marble?

Answer by rothauserc(4718)

probability (P) of picking one red marble, and one green marble is
:
there are 11C2 ways to pick two marbles out of 11 without replacement, where 11C2 = 11!/(2! * (11-2)!) = 55 ways
:
then we have
:
(4C1 * 7C1) / 55 = 28/55 = 0.5091
: